# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v3

@@ 1.2 Model Inference and Global Embeddings Maps
Following pretraining, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10 m pixel globally. This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders. To ensure that our maps are gap-free, random sampling with replacement is used to generate augmentations from d-pixels that have fewer than 40 non-cloudy observations. The primary output of this inference stage is a set of annual, global, 10 m resolution TESSERA embedding maps. These maps are designed as readily usable multichannel geospatial data layers. This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself (see Appendix A.8). We are generating global embeddings for 2024 (approx. 250 TB) and will progressively extend coverage backward year by year until 2017. Visualizing the first three dimensions of the embeddings as R,G, and B channels reveals rich, detailed textures corresponding to diverse land cover and land use patterns. These visualizations highlight the model’s ability to capture information across multiple scales, from individual agricultural fields to entire continents (Fig. 2).
N
@@
2 Downstream Tasks
The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks. These include pixel-wise classification, such as for crop type mapping, pixel-wise regression, such as for estimating canopy height and carbon density of forests, and patch-based dense predictions, such as for semantic segmentation. (See Appendix D and Supplementary Fig. 12 for the general downstream application workflow.) In these evaluations, the pre-trained TESSERA encoder, with frozen weights, acts as a feature extractor (although, given adequate computing resources, fine-tuning the model is also possible), converting d-pixels from a region of interest to embeddings. Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these embeddings, acting as explanatory variables in regression modelling. This approach is computation- and labelled-data efficient: it typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state-of-the-art.
2.1 Crop Type Classification
Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector [177]. The classification of crop types informs vital agricultural decisions by allowing accurate estimates of crop area and production, and facilitating low-cost monitoring of diseases and pests in vast regions [61]. However, this task is considerably more complex than many other challenges in the classification of land cover, as it requires discerning subtle spectral and temporal variations that differentiate different crop classes [61]. In the past, classifiers such as Random Forest (RF) applied to hand-crafted composite time series data have established the standard [61, 113]. Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific feature extraction and training, offer generalizable representations that can be applied universally [7, 80]. Here, we demonstrate that TESSERA’s precomputed embeddings set a new state-of-the-art.
To evaluate performance, we used the INVEKOS Austrian crop dataset [3] from the 2021–2022 growing season. This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets) [3]. We compared the performance of a small classifier neural network trained using TESSERA embeddings with three key state-of-the-art alternatives: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model [151], and (3) Google Satellite Embedding (GSE) [57]. For the foundation models, we used the same small neural network to classify their respective embeddings to allow a direct comparison. We evaluated performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labelled data to one-shot learning. We also evaluated performance on a patch-based semantic segmentation task against several other foundation models. Details on data processing, model architectures, and training protocols are provided in Appendix E.
These performance gains are due to the high quality of the learned embeddings. A UMAP analysis [102] of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from PRESTO and GSE (Fig. 3d). This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.
@@ 2.1 Crop Type Classification
2.2 Canopy height estimation
Canopy height is a key structural attribute of forests, linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks such as REDD+ [105, 50]. It is correlated with ecosystem function [104], forest age, and successional stage, while also serving as a proxy for habitat quality, an important determinant of biodiversity patterns and species richness [38, 26, 138]. Accurate, scalable canopy height models (CHMs) can enable detection of forest degradation and disturbance.
Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability. Airborne Laser Scanning (ALS), where laser-based LiDAR instruments are carried in an aircraft flying over a target region, creates highly accurate benchmark datasets, but is necessarily restricted in geographical range [23]. GEDI, a spaceborne LiDAR system, can provide accurate point measurements of canopy height to train models to map canopy height from optical and radar data; however, it does not provide wall-to-wall maps, does not sample high latitudes [42], and does not produce unreliable estimates in mountainous terrain [53]. Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height [132]. Optical data, while widely available, may lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [160, 96] - yet none of the three most widely used global CHM models currently implement sensor fusion [84, 120, 150]. Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500> 500 features in Potapov et al. [120]), which may generalize poorly across ecosystems. These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [83], and inconsistencies in input quality. As a result, significant discrepancies remain between global CHMs and ALS benchmarks, and current products underestimate local variability and structural extremes [106]. Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration. These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [107, 115].
We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo [75] with openly available ground-truth data [34]. To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation. A 30-million-parameter UNet was trained using 32 $\times$ 32 pixel patches of the embeddings as input. To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set. The accuracy metrics for the foundation models are reported as the average across the 12 runs. TESSERA achieved a coefficiento of determination ($R^{2}$) of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62- 0.62 m.
We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig. 4). First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al. [120], the ETH map by Lang et al. [84], and the Meta map by Tolan et al. [150]. Although local regression models, such as ours, generally outperform global products in regional evaluations [128], the pronounced saturation of these global models at much lower canopy heights–combined with RMSEs that exceed their reported global averages [120, 84, 150]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
We also compared our results with regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al. [85], which used deep convolutional neural networks trained in Sentinel-2 time series and GEDI LiDAR to estimate the regional canopy height and biomass. Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO [151] and Google Satellite Embedding (GSE) [57], using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models. Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
@@ 2.2 Canopy height estimation
Figure 4: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation. Predicted canopy height (y-axis) is compared against airborne LiDAR reference data (x-axis) across multiple models and input datasets in Danum Valley, Borneo. Top row: Predictions from models trained using different input representations—TESSERA, PRESTO, Google Satellite Embedding (GSE), and raw Sentinel-2 composites—using the same model architecture and training procedure. TESSERA achieves the highest coefficient of determination and the lowest RMSE (8.88±0.988.88$\pm$ 0.988.88 ± 0.98 m), indicating superior performance in structurally complex, tall tropical canopies. Bottom row: Comparisons to existing global and regional canopy height maps, including products from ETH, GLAD, Meta, and Lang et al.’s regional CHM. All four products show substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14$ m), with three exhibiting negative correlations. These models saturate at lower canopy heights and show large bias and RMSE, highlighting the limitations of global products when applied to high-biomass tropical forests without local calibration. A 1:1 reference line (blue) is shown in each panel for comparison.
2.3 Above-ground forest biomass estimation
Forests play a central role in the global carbon cycle [67]. This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery [46]. A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions. Thus, there has been a long-standing effort to link field measurements to global satellite observations [194].
Recently, with the creation of the BioMassters benchmark dataset [110] (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice  [5]. We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98. In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT [68] and GSE [57].
First, a summary of the BioMassters dataset. This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations [126, 127] informed by extensive field plots. Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel. The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values (Fig. 5b) that would not appear in coarser resolution data. To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
We compared a UNet-based regression model [129] trained on TESSERA embeddings with SpectralGPT embeddings [68] combined with a lightweight spatial-temporal encoder [133] for temporal aggregation and a UPerNet [181] regression head, because this combination has been found in the Pangaea benchmark [100] to outperform other foundation models [110, 100]. Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase. Thus, we adopted the Pangaea implementation of their model [100], which uses preprocessed satellite imagery from the BioMassters dataset. We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline. This model uses a UNet with a temporal attention encoder [110]. Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input. However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable. Finally, we compared our work with a UNet-based regressor trained on GSE embeddings.
@@ 2.3 Above-ground forest biomass estimation
2.4 Burned Area Detection
The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change [6, 72, 116]. Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends [11], predicting wildfire occurrence [30], assessing ecosystem recovery [188], and establishing effective fire management strategies [64].
Remote sensing is widely used to detect burned areas and assess severity regionally and globally [82]. Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [82, 118]. These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [82]. Most existing methods use optical sensors [82], which can detect fire-specific spectral signatures but are susceptible to cloud cover [65], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [66]. SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [82, 145, 156]. Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness [82, 31]. Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [145, 156, 130]. TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig. 6).
@@ 2.4 Burned Area Detection
Figure 6: Fire disturbance detection using TESSERA embeddings across two burned regions in California. a, UMAP projection of 2020 TESSERA embeddings reveals clear structure separating unburned pixels and two temporally distinct fire events. A continuous gradient within the August fire aligns with MTBS burn severity classes, suggesting that TESSERA captures both timing and magnitude of disturbance. b, Burned vs. unburned classification performance using random forest classifiers trained on different input representations. TESSERA achieves high F1-scores with minimal training data and performs comparably to a task-specific composite baseline, while significantly outperforming Google Satellite Embeddings (GSE). c, Burn classification maps generated from models trained on 1% of labelled data demonstrate strong agreement with MTBS perimeters across both test areas. As MTBS is itself derived from remote sensing imagery, discrepancies may reflect methodological differences rather than prediction error.
We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Figure 6). Our analysis focused on two randomly selected 15$\times$15 km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset [4]. UMAP projections [102] of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters. In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Figure 6a). To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels. We compared three input representations: (1) TESSERA embeddings, (2) Google Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post-fire Sentinel-2 composites. The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Figure 6 b). Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency. These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
2.5 Stocking indices in voluntary carbon markets
The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere [109]. To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas. This is a costly and inefficient undertaking if it is done only by field sampling.
To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [157]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB. Upon acceptance, project developers can use the index to assess the performance of the project. Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM. To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al. [14]. Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see  [1] for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil. This rescaling ensures a fair comparison between the various products.

### v4

@@ Abstract
Following pre-training, the TESSERA dual encoder (with frozen weights and excluding the projector) is used to generate an annual 128-dimensional embedding for every 10 m pixel globally. This involves making inferences from a sample of length 40 from the annual Sentinel-1 and Sentinel-2 time series for each pixel through the trained encoders. To ensure that our maps are gap-free, random sampling with replacement is used to generate augmentations from d-pixels that have fewer than 40 non-cloudy observations. The primary output of this inference stage is a set of annual, global, 10 m resolution TESSERA embedding maps. These maps are designed as readily usable multichannel geospatial data layers. This “Embeddings-as-Data” approach significantly lowers the barrier to entry for end-users, as these rich, precomputed features can be directly ingested by downstream models without the need for raw satellite data processing or running the TESSERA model itself (see Supplementary Text ). We are currently generating global embeddings for 2024 (approx. 250 TB) and will progressively extend coverage backward year by year until 2017. Visualizing the first three dimensions of the embeddings as R,G, and B channels reveals rich, detailed textures corresponding to diverse land cover and land use patterns. These visualizations highlight the model’s ability to capture information across multiple scales, from individual agricultural fields to entire continents (Fig. 2).
Downstream Tasks
The effectiveness and generalizability of the TESSERA embeddings are evaluated on a diverse range of downstream remote sensing tasks. These include pixel-wise classification, such as for crop type mapping, pixel-wise regression, such as for estimating canopy height and carbon density of forests, and patch-based dense predictions, such as for semantic segmentation. (See Supplementary fig. S5 for the general downstream application workflow.) In these evaluations, the pre-trained TESSERA encoder, with frozen weights, acts as a feature extractor (although, given adequate computing resources, fine-tuning the model is also possible), converting d-pixels from a region of interest to embeddings. Lightweight, task-specific model heads (e.g., shallow MLPs or UNet-style decoders) are then trained using these embeddings, acting as explanatory variables in regression modelling. This approach is computation- and labelled-data efficient: it typically requires substantially less labelled data and computational resources compared to training deep models from scratch, yet achieves or surpasses the state-of-the-art.
Crop Type Classification
Agricultural monitoring through EO is a cornerstone of managing global food security, providing critical data to stakeholders in the agricultural sector  (?). The classification of crop types informs vital agricultural decisions by allowing accurate estimates of crop area and production, and facilitating low-cost monitoring of diseases and pests in vast regions  (?). However, this task is considerably more complex than many other challenges in the classification of land cover, as it requires discerning subtle spectral and temporal variations that differentiate different crop classes  (?). In the past, classifiers such as Random Forest (RF) applied to hand-crafted composite time series data have established the standard  (?, ?). Recent advances have shifted towards foundation models, which, unlike bespoke machine learning models that require region-specific feature extraction and training, offer generalizable representations that can be applied universally  (?, ?). Here, we demonstrate that TESSERA’s precomputed embeddings set a new state-of-the-art.
To evaluate performance, we used the INVEKOS Austrian crop dataset  (?) from the 2021–2022 growing season. This is an extensive data set that covers 1850 km2 east of Vienna with 154 different crop types, which we consolidated into 17 classes based on crop similarity (e.g., merging feed beet with sugar beets)  (?). We compared the performance of a small classifier neural network trained using TESSERA embeddings with three key state-of-the-art alternatives: (1) an RF classifier trained on engineered temporal features from raw satellite data, (2) the PRESTO foundation model  (?), and (3) AlphaEarth Google Satellite Embedding (GSE)  (?). For the foundation models, we used the same small neural network to classify their respective embeddings to allow a direct comparison. We evaluated performance using multiple metrics, including F1 scores, under varying data availability scenarios, from 30% of labelled data to one-shot learning. We also evaluated performance on a patch-based semantic segmentation task against several other foundation models. Details on data processing, model architectures, and training protocols are provided in the Supplementary Text.
These performance gains are due to the high quality of the learned embeddings. A UMAP analysis  (?) of the embedding space reveals that TESSERA embeddings form tighter, more semantically coherent clusters by crop class compared to those from PRESTO and GSE (Fig. 3d). This improved separability highlights the strength of TESSERA as a foundational tool for agricultural applications.
Canopy Height Estimation
Canopy height is a key structural attribute of forests, linked to above-ground biomass and carbon stocks, and therefore important for climate mitigation efforts and carbon accounting frameworks such as REDD+  (?, ?). It is correlated with ecosystem function  (?), forest age, and successional stage, while also serving as a proxy for habitat quality, an important determinant of biodiversity patterns and species richness  (?, ?, ?). Accurate, scalable canopy height models (CHMs) can enable detection of forest degradation and disturbance.
Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability. Airborne Laser Scanning (ALS), where laser-based LiDAR instruments are carried in an aircraft flying over a target region, creates highly accurate benchmark datasets, but is necessarily restricted in geographical range  (?). GEDI, a spaceborne LiDAR system, can provide accurate point measurements of canopy height to train models to map canopy height from optical and radar data; however, it does not provide wall-to-wall maps, does not sample high latitudes  (?), and does not produce unreliable estimates in mountainous terrain  (?). Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height  (?). Optical data, while widely available, may lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors   (?, ?) - yet none of the three most widely used global CHM models currently implement sensor fusion  (?, ?, ?). Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500 features in Potapov et al.  (?)), which may generalize poorly across ecosystems. These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation  (?), and inconsistencies in input quality. As a result, significant discrepancies remain between global CHMs and ALS benchmarks, and current products underestimate local variability and structural extremes  (?). Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration. These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling  (?, ?).
We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo  (?) with openly available ground-truth data  (?). To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation. A 30-million-parameter UNet was trained using 32 $\times$ 32 pixel patches of the embeddings as input. To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set. The accuracy metrics for the foundation models are reported as the average across the 12 runs. TESSERA achieved a coefficiento of determination ($R^{2}$) of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62 m.
We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig. 4). First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al.  (?), the ETH map by Lang et al.  (?), and the Meta map by Tolan et al.  (?). Although local regression models, such as ours, generally outperform global products in regional evaluations  (?), the pronounced saturation of these global models at much lower canopy heights–combined with RMSEs that exceed their reported global averages  (?, ?, ?)—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
We also compared our results with regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al.  (?), which used deep convolutional neural networks trained in Sentinel-2 time series and GEDI LiDAR to estimate the regional canopy height and biomass. Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO  (?) and Google Satellite Embedding (GSE)  (?), using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models. Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
Above-ground Forest Biomass Estimation
Forests play a central role in the global carbon cycle  (?). This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery  (?). A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions. Thus, there has been a long-standing effort to link field measurements to global satellite observations  (?).
Recently, with the creation of the BioMassters benchmark dataset  (?) (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice   (?). We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98. In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT  (?) and GSE  (?).
First, a summary of the BioMassters dataset. This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations  (?, ?) informed by extensive field plots. Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel. The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values (Fig. 5b) that would not appear in coarser resolution data. To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
We compared a UNet-based regression model  (?) trained on TESSERA embeddings with SpectralGPT embeddings  (?) combined with a lightweight spatial-temporal encoder  (?) for temporal aggregation and a UPerNet  (?) regression head, because this combination has been found in the Pangaea benchmark  (?) to outperform other foundation models  (?, ?). Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase. Thus, we adopted the Pangaea implementation of their model  (?), which uses preprocessed satellite imagery from the BioMassters dataset. We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline. This model uses a UNet with a temporal attention encoder  (?). Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input. However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable. Finally, we compared our work with a UNet-based regressor trained on GSE embeddings.
Burned Area Detection
The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change  (?, ?, ?). Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends  (?), predicting wildfire occurrence  (?), assessing ecosystem recovery  (?), and establishing effective fire management strategies  (?).
Remote sensing is widely used to detect burned areas and assess severity regionally and globally  (?). Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input  (?, ?). These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals  (?). Most existing methods use optical sensors  (?), which can detect fire-specific spectral signatures but are susceptible to cloud cover  (?), particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens  (?). SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover  (?, ?, ?). Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness  (?, ?). Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products  (?, ?, ?). TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig. 6).
We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Figure 6). Our analysis focused on two randomly selected $\times15$ km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset  (?). UMAP projections  (?) of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters. In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Figure 6a). To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels. We compared three input representations: (1) TESSERA embeddings, (2) Google Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post-fire Sentinel-2 composites. The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Figure 6 b). Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency. These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
Stocking Indices in Voluntary Carbon Markets
The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere  (?). To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas. This is a costly and inefficient undertaking if it is done only by field sampling.
To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra  (?)) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB. Upon acceptance, project developers can use the index to assess the performance of the project. Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM. To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al.  (?). Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see   (?) for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil. This rescaling ensures a fair comparison between the various products.
TESSERA: a Global Open-Source Foundation Model
Figure 1: Data processing pipeline. a, Overview. Time-series data from Sentinel-2 MSI and Sentinel-1 SAR observations are extracted for each pixel location to form mode-specific ‘d-pixels’ ( vertical rectangles; cloud-corrupted rows in gray). Two augmentations are created from each d-pixel as cloud-free random samples from the corresponding d-pixel. Augmentations are processed through a dual-branch architecture with weights shared between the branches. The two resulting 128-dimensional embeddings are fused by a multilayer perceptron (MLP) to create the final output embeddings. During training (but not inferencing) the fused features are expanded to 16,384 dimensions by a large projector network. A modified Barlow Twins loss (a linear combination of $\mathcal{L}_{BT}$ and $\mathcal{L}_{MIX}$) is calculated from the cross-correlation matrix of projected features, forcing the model to learn dimensionally decorrelated embeddings that are invariant to masked days in the Sentinel-2 d-pixel. b, Detailed structure of a single encoder branch. Each d-pixel is embedded using a single-layer MLP (not shown) and a Day-of-Year-based temporal encoding is added. The resulting vector is processed by an 4-block Transformer Encoder and then a Gated Recurrent Unit (GRU) pooling layer; its final hidden state, which summarizes the entire time series, serves as the single 128-dimensional embedding for that modality, which is then passed to the fusion MLP shown in panel a.
@@
Figure 4: TESSERA outperforms global and regional models as well as alternative pixel-based foundation models in canopy height estimation. Predicted canopy height (y-axis) is compared against airborne LiDAR reference data (x-axis) across multiple models and input datasets in Danum Valley, Borneo. Top row: Predictions from models trained using different input representations—TESSERA, PRESTO, Google Satellite Embedding (GSE), and raw Sentinel-2 composites—using the same model architecture and training procedure. TESSERA achieves the highest coefficient of determination and the lowest RMSE ($8.88\pm 0.98$ m), indicating superior performance in structurally complex, tall tropical canopies. Bottom row: Comparisons to existing global and regional canopy height maps, including products from ETH, GLAD, Meta, and Lang et al.’s regional CHM. All four products show substantially lower predictive performance ($R^{2}<0.05$, $RMSE>14$ m), with three exhibiting negative correlations. These models saturate at lower canopy heights and show large bias and RMSE, highlighting the limitations of global products when applied to high-biomass tropical forests without local calibration. A 1:1 reference line (blue) is shown in each panel for comparison.
@@
Figure 6: Fire disturbance detection using TESSERA embeddings across two burned regions in California. a, UMAP projection of $20$ TESSERA embeddings reveals clear structure separating unburned pixels and two temporally distinct fire events. A continuous gradient within the August fire aligns with MTBS burn severity classes, suggesting that TESSERA captures both timing and magnitude of disturbance. b, Burned vs. unburned classification performance using random forest classifiers trained on different input representations. TESSERA achieves high F1-scores with minimal training data and performs comparably to a task-specific composite baseline, while significantly outperforming Google Satellite Embeddings (GSE). c, Burn classification maps generated from models trained on 1% of labelled data demonstrate strong agreement with MTBS perimeters across both test areas. As MTBS is itself derived from remote sensing imagery, discrepancies may reflect methodological differences rather than prediction error.